OpenStudy (meilendurcer):

Write a percent equation that can be used to solve the following problem. 25 is what percent of 350?

OpenStudy (anonymous):

3. Write a percent equation that can be used to solve the following problem: 25 is what percent of 350? a. n = 25 × 350 b. 350 = p × 25 c. 0.25 = p × 350 d. none of these the answer is: d. none of these

OpenStudy (meilendurcer):

oh you do this school too?

OpenStudy (meilendurcer):

thanks

OpenStudy (anonymous):

your welcome @meilendurcer

OpenStudy (anonymous):

\[\frac{ x}{ 100 }*350=25\\ \frac{ x}{ \cancel{10}*10 }*\cancel{10}*35=25\\ x= 87.5\]